OpenStudy (anonymous):

show that if k is a positive constant, then the area between the x-axis and one arch of the curve y=sin kx is 2/k? need help :(

OpenStudy (anonymous):

Allowed to use Calculus?

OpenStudy (anonymous):

\[ \huge \int_0^{\frac{\pi}{k}}\sin(kx)dx \]